Given that you know one gas in a mixture of gases has a pressure of .5 atm, and has a mol fraction of .18, what volume will the mixture of gases occupy if the temperature is 390 kelvin, the total number of moles of all the gases combined are 20?
1
Expert's answer
2015-11-17T08:57:49-0500
pV = vRT
V = vRT/p
R = 0.082 L atm/K mol

V = 20 * 0.082 * 390 / 0.5 = 1279.2 L